OpenStudy (anonymous):

Find the area of the triangle with vertices (3,5), (2,-7), (-4,-6). Prove by graphing

OpenStudy (anonymous):

am I right? that after finding the distance between two points... all I have to do is find its area.

OpenStudy (anonymous):

A = 1/2 (base) * (height). So as long as the two sides give the base and height, you should be able to figure it out.

OpenStudy (anonymous):

You can use the A=1/2(base)*(height) only if you find the sides to be Pythagorean triplets (third side = sqrt(sum of squares of other two sides)). Otherwise use the formula A = sqrt[s*(s-a)*(s-b)*(s-c)] where s=(a+b+c)/2 and a, b, c are the sides of the triangle. The second formula is called Hero's formula :D

OpenStudy (anonymous):

There is a formula for this if u are allowed to use it. Label the points A,B,C Area is absolute value of Ax(By-Cy)+ Bx(Cy-Ay) + Cx(Ay-By) all over 2 (Ax,Ay coordinates of A, etc)

OpenStudy (anonymous):

thank you estudier